OpenStudy (anonymous):

how do you find the inverse of log12(x)? i need lots of help THANKS :)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

also, how would you find the domain or range of it? i am very confused wiht logs.

OpenStudy (anonymous):

inverse is \[f(x)=12^x\] domain of log is \[(0,\infty)\] range is \[(-\infty, \infty)\]

OpenStudy (anonymous):

thank you very much!!!!